Session: May 1997 (session 2)
Frequency: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (T. Venturi at Tue Jun 10 20:47:17 METDST 1997) Success
Experiment ran smoothly, without apparent problems. Clear but very humid weather

Noto (C.Trigilio at Wed Jun 11 12:40:00 METDST 1997) Success
No problem reported

Torun (K.Borkowski at Wed Jun 11 14:45:56 METDST 1997) Success

Urumqi (Li Suqin at Thu Jun 12 07:27:52 METDST 1997) Success
It is fine day , the colck offset =8.2+-0.5us. No problem found.

Jodrell (Alastair Gunn at Mon Jun 16 11:34:50 METDST 1997) Success
No known problems.

Effelsberg (dgraham at Mon Jun 16 14:35:20 METDST 1997) Minor failures
late start at 13:08

Simeiz (A.Volvatch/LIG at Wed Jun 18 09:06:14 METDST 1997) Success
No problem reported, tapes shipped on 17.06.97.

Onsala (John Conway at Wed Jun 18 13:42:22 METDST 1997) Minor failures
Lost first few hours of expt yp 08:51 due to operatore error, shedule was started late.

HartRAO (Jonathan Quick at Mon Jul 7 11:22:57 METDST 1997) Success
No problems

Medicina (Fredrik T Rantakyro at Mon Jul 21 11:21:11 METDST 1997) Minor failures
The ANTAB file for the experiment can be found on VLBEER in the MAY97 directory with the name EH2B.ANTAB. The file is ready to be used with the AIPS tasks ANTAB/APCAL. Read the ANTAB file for any specific comments regarding this experiment.